Can the work be shown in excel using formulas?


You deposit $10,000 in an account that earns 5.5% compounded continuously and your friend deposits $10,000 in an account that earns 5.5% annually.

a. How much more will you have in the account in 10 years?
b. How much more interest did you earn in the 10 years?
Answer #1

The compounded or Future value (A) in case of annually compounded and continuously compounded rate r over a period of t years on a principal amount P is given by

A = P* (1+r)^t ------ in case of Annually compounded and

A = P*exp (r*t) -------- in case of continuously compounded

Amount in my friend's account = 10000 * (1+0.055)^10 = $17081.44

Amount in my account = 10000 * exp(0.055*10) = $17332.53

a) After 10 years, I will have = 17332.53 - 17081.44 = $251.09 more in my account as compared to my friend

b) Difference in Interest = ($17332.53-$10000) + ($17081.44-$10000) =$251.09

I earned $251.09 more interest as compared to my friend

Shown in excel below

Principal Rate Final Amount after 10 years
My Friend 10000 5.5% annually compounded 17081.44
Me 10000 5.5% continuously compounded 17332.53
Difference in Amount 251.0856